OpenStudy (anonymous):

Alonzo drove to the mountains last weekend. There was heavy traffic on the way there, and the trip took 10 hours. When Alonzo drove home, there was no traffic and the trip only took 7 hours. If his average rate was 18 miles per hour faster on the trip home, how far away does Alonzo live from the mountains? Do not do any rounding. I know I'm setting it up wrong because I keep getting a negative answer. Here is my work, Rate Time Distance Trip 1 x 10 x Trip 2 x+18 7 x So, I did x=10x+7(x+18)

OpenStudy (anonymous):

roofies lol

OpenStudy (phi):

you are close. But you must use a different variable name for distance, because x cannot be both speed and distance.

OpenStudy (phi):

Rate * Time = Distance Trip 1 x * 10 = y Trip 2 (x+18) * 7 = y

OpenStudy (anonymous):

So y=10x+7(x+18)

OpenStudy (phi):

the distance is the same for both trips, so we can set both to the same thing (y) ** So y=10x+7(x+18) *** no. you have Trip 1 x * 10 = y Trip 2 (x+18) * 7 = y

OpenStudy (phi):

you might notice that y equals y and use that idea

OpenStudy (anonymous):

ohh, so two different equations. 10x=7x+126 3x=126 x=42

OpenStudy (phi):

yes.